# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v1

@@ I Introduction
is neither zero nor infinity in the spatial region in which they are to be used. Defining the position function just means giving $x$, $y$, and $z$ as functions of $(\psi,\theta,\varphi)$. The appendix to (Boozer, 2004) explains how arbitrary coordinate systems, such as $\vec{x}(\psi,\theta,\varphi)$, are used in calculations.
@@
The reason coordinate systems are important to a discussion of the properties of magnetic field lines, is the trajectories are given by a Hamiltonian of the $H(p,q,t)$ form. The Hamiltonian, its canonical momentum $p$ and canonical coordinate $q$, and its canonical time $t$ have a subtle relationship to physical quantities. Section II shows the Hamiltonian is the poloidal flux $\psi_{p}(\psi_{t},\theta,\varphi,t)$, the canonical momentum is the toroidal magnetic flux $\psi_{t}$, the canonical coordinate is the poloidal angle $\theta$, and the canonical time the toroidal angle $\varphi$, Figure 1. The actual clock time $t$ is just a parameter in the Hamiltonian.
@@
(Boozer, 1983) showed the topological properties of magnetic field lines at each instant in time are given by $\psi_{p}(\psi_{t},\theta,\varphi,t)$. The position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ is needed to plot the field lines in ordinary space with $\partial\vec{x}/\partial t$ giving the velocity of $(\psi_{t},\theta,\varphi)$ through space, which is the field line velocity when field line topology is not changing, which implies $\psi_{p}$ can be taken to be independent of time.
@@
When a particular line $\vec{x}_{0}(\ell)$ remains within a bounded region of space, it has three possibilities as $\ell\rightarrow\infty$: It can close on itself after a distance $\ell_{0}$. It can cover a surface coming arbitrarily close to every point on a surface without ever going through the same point twice. It can cover a volume coming arbitrarily close to every point in volume of space without ever going through the same point twice. All three types of trajectories are illustrated by the Standard Map, Figure (2). The iteration number represents the distance $\ell$. The periodicity of both coordinates of the Standard Map makes it non-trivial to translate them into the $(\psi_{t},\theta)$ coordinates of the magnetic field line Hamiltonian $\psi_{p}$. (Boozer & Punjabi, 2018) in their Appendix B have obtained an explicit formula for a $\psi_{p}(\psi_{t},\theta,\varphi)$ that represents the qualitative features of stellarator equilibria including the non-resonant divertor that (Strumberger, 1992) found naturally arises at the edge optimized stellarators. (Bader et al., 2018) found that the locations where the field lines strike the walls in non-resonant divertors are remarkably robust to changes in the magnetic field
@@
Figure 2: Trajectories given by iterations of the Standard Map are illustrated at k=0.975. Black trajectories show island chains. The red region is a single trajectory iterated $10^{6}$ times but still escapes from that region after approximately $10^{9}$ iterations. The blue and the yellow regions are similar single-trajectory chaotic regions from which the trajectory escapes after a very large number of iterations. The $n+1$ iterate of the standard map is $\alpha_{n+1}=\alpha_{n}+\beta_{n+1}$ and $\beta_{n+1}=\beta_{n}-(k/2\pi)\sin(2\pi\alpha_{n})$. Both $\alpha$, the horizontal axis, and $\beta$, the vertical axis, are periodic with a period of unity. This was Figure 1 in J. D. Meiss, Thirty years of turnstiles and transport, Chaos 25, 097602 (2015).
@@
The behavior of the magnetic field lines that are infinitesimally separated from an arbitrary line $\vec{x}_{0}(\ell)$ can be studied without approximation by a particularly simple Hamiltonian (Boozer, 2012). The study of neighboring lines, which is reviewed in Appendix A, showed: (1) It is possible for field lines infinitesimally separated from a given line $\vec{x}_{0}(\ell)$, of any of the three types, to have separations that increase or decrease exponentially with $\ell$, Appendix A.1. There must be an equal number of neighboring lines that diverge from and approach the line $\vec{x}_{0}(\ell)$ due to $\vec{\nabla}\cdot\vec{B}=0$. A circular flux tube that is defined about the line $\vec{x}_{0}(\ell)$ at $\ell=0$ distorts into an ellipse with the ratio of the major to the minor axis equal to $e^{2\sigma_{max}(\ell)}$. (2) When the magnetic field depends on only two spatial coordinates, a single field line cannot cover a volume, as in the red area of Figure 2. Three spatial coordinates are required. Two-coordinate magnetic fields, as in axisymmetry, can have X-points which are specific field lines from which neighboring lines exponentially approach and separate. However, the volume in which lines have at least a certain exponentiation is itself exponentially small. (3) Curl-free magnetic fields can exhibit field lines with an exponential separation throughout a volume. An ideal magnetic evolution does not conserve the exponentiation $\sigma_{max}(\ell)$. In other words, the exponentiation can be a property of the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ with $\psi_{p}$ an unchanging function of its canonical coordinates. The current density required for an ideal evolution to produce an exponentiation $\sigma_{max}$ is proportional to $\sigma_{max}$, Appendix A.3. As will be discussed in Section VI, the time required for a large scale reconnection to occur is determined by a timescale set by the ideal evolution times $\sigma_{r}$. The required exponentiation $\sigma_{r}$ is the ratio of timescale for non-ideal diffusive effects, such as resistivity, to a timescale determined by the ideal evolution, Section V.
@@
The tendency to construct two dimensional models has led to the prejudice that exponential separation of magnetic field lines throughout a volume is peculiar. Actually all magnetic fields that depend on all three spatial coordinates tend to be chaotic, which is illustrated by the design of a curl-free magnetic fields for a stellarators. Obtaining nested magnetic surfaces, as required for a stellarator, rather than having the field lines fill a volume requires careful optimization. Even a small perturbation to an optimized curl-free stellarator can destroy the surfaces. Nevertheless, axisymmetric tokamaks are even more sensitive to external field perturbations due to their amplification by coupling to marginally stable current-driven kinks (Park et al., 2010).
@@
The concept of magnetic field line chaos and its definition are sufficiently subtle that Section IV is largely devoted to the topic.
@@
When the magnetic field lines lie in toroidal magnetic surfaces, the poloidal angle $\theta$ can be chosen so the poloidal flux is a function of only the toroidal flux and time, $\psi_{p}(\psi_{t},t)$. At each instant in time, the trajectories of field lines are then given by $\psi_{t}$ and $\theta_{0}$ constant with $\theta=\theta_{0}+\iota(\psi)\varphi$, where the rotational transform $\iota\equiv(\partial\psi_{p}/\partial\psi)_{t}$. In tokamaks, it is customary to use the safety factor $q\equiv 1/\iota$ instead of the rotational transform.
@@
On a rational surface the rotational transform is a rational number, the ratio of two integers $\iota=N/M$. The field lines on that surface close after $N$ poloidal transits and $M$ toroidal transits. This identification with poloidal and toroidal transits may seem backwards from the standard convention of using $m$ as a poloidal and $n$ as a toroidal mode number as in $\cos(m\theta-n\varphi)$. But, when $\iota=N/M$, then along a field line $\cos(m\theta-n\varphi)=\cos\Big(m\theta_{0}+(mN-n)\varphi)\theta-n\varphi)\Big)$, which has no variation along the line since it equals $\cos(m\theta_{0})$ when $m=M$ and $n=N$.
@@
(Boozer, 2004) showed the time derivative of the poloidal flux is the loop voltage,
$\frac{\partial\psi_{p}(\psi,t)}{\partial t}=V_{\ell}.$ (2)
The loop voltage is defined as an integral along a field line at an instant of time,
@@
When $\vec{\nabla}V_{\ell}=0$, the topology of the magnetic field lines cannot change, and their evolution is considered ideal. Nevertheless, the poloidal flux has an additive position-independent function of time when $\vec{\nabla}V_{\ell}=0$ but $V_{\ell}\neq 0$.
@@
Section III explains a perturbation that breaks magnetic surfaces amd forms chains of magnetic islands at the rational surfaces that resonate with the rotational transform $\iota\equiv\partial\psi_{p}/\partial\partial_{t}$. A resonant perturbation to $\psi_{p}$ has a toroidal mode number $n$ and a poloidal mode number $m$ with $n/m=\iota$. The width of these islands is proportional to the square root of the resonant perturbations. Non-resonant perturbations distort magnetic surfaces but do not break them. Resonant perturbations that create islands cannot be ideal, they must change the form of the poloidal flux, but the position function can be held fixed.
@@
Non-resonant perturbations can be taken to change the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ while the poloidal flux can be held fixed using a canonical transformation or what is mathematically equivalent an electric potential as shown in the appendix to (Boozer, 2004).
If the plasma evolution were perfectly ideal, a resonant perturbation would affect only the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$. But, a resonant ideal perturbation forces the magnetic field to have a delta-function current density at the resonant rational surface $\iota=n/m$. This current causes distortions to the neighboring surfaces producing magnetic perturbations with all harmonics $k$ of the resonant perturbation $(kn)/(km)=\iota$ with $k$ any integer (Huang et al., 2022). In a torus, the toroidicity produces magnetic perturbations with poloidal harmonics of strength $\epsilon^{k}$ at $m\pm k$ with $\epsilon$ the inverse aspect ratio of the torus. Any rational number can be produced by the harmonics of $n/m$ together with the harmonics of $m$ produced by toroidicity. The more ideal the perturbation, the more contorted the magnetic surfaces become, which leads to their rapid breaking by an arbitrarily small non-ideal effect, such as resistivity (Boozer, 2022). This is discussed in Section VI.
@@
(Chirikov, 1960) gave the approximate condition for the creation of large chaotic volumes: the overlap of the islands from different resonant rational surfaces. The actual formation of chaotic regions is more complicated than just the overlap of island chains. (MacKay, Meiss & Percival, 1984) showed that chaotic regions are formed by cantori and the turnstiles that penetrate them.
@@
The last magnetic surface to break between two large magnetic island chains is the least rational surface between the resonant surfaces associated with those chains, Section V. This is quantified by the magnitude of $m$ required for the transform on this surface to satisfy $\iota(\psi_{t})=n/m+\epsilon_{r}$ with $n$ a toroidal mode number and $m$ a poloidal mode number as $|\epsilon_{r}|\rightarrow 0$. As the last magnetic surface breaks, it is called a cantorus. A cantorus is like an irrational magnetic surface but develops holes. These holes are places where $\vec{B}\cdot\hat{n}\neq 0$ with $\hat{n}$ the unit normal to the cantorus. The holes define highly collimated tubes of magnetic flux and must come as an inward and an outward pair due to $\vec{\nabla}\cdot\vec{B}=0$. For this reason they are called turnstiles.
@@
The highly collimated flux tubes of turnstiles are responsible for the severe damage that can be caused by runaway electrons following the lines in these tubes to localized positions on the chamber walls (Boozer & Pujabi, 2016). This collimation also responsible for the formation of a natural non-resonant divertor at the edge of a stellarator (Strumberger, 1992) and (Boozer & Punjabi, 2018).
@@
Section VII gives a few examples of the application if the concepts of magnetic field line chaos, cantori, and turnstiles to toroidal fusion plasmas but also expresses concern about the future of innovation within U. S. fusion program.
II Representation of a magnetic field and its field lines
@@
(Boozer, 1983) showed that the magnetic field in a toroidal region can always be written as
@@
Equation (4) follows from writing the vector potential $\vec{A}$, where $\vec{B}=\vec{\nabla}\times\vec{A}$, using three coordinates $(r,\theta,\varphi)$. Then, $\vec{A}=A_{r}\vec{\nabla}r+A_{\theta}\vec{\nabla}\theta+A_{r}\vec{\nabla}\varphi+\vec{\nabla}g$. The gauge can be chosen $\partial g(r,\theta,\varphi)/\partial r=-A_{r}$. Let $\psi_{t}/2\pi\equiv A_{\theta}+\partial g/\partial\theta$, and $\psi_{p}/2\pi\equiv-(A_{\varphi}+\partial g/\partial\varphi)$, so $2\pi\vec{A}=\psi_{t}\vec{\nabla}\theta-\psi_{p}\vec{\nabla}\varphi$, which has Equation (4) as its curl.
In general the poloidal flux depends on all three spatial coordinates and time, $\psi_{p}(\psi_{t},\theta,\varphi,t)$. (Boozer, 1983) showed the trajectories of the magnetic field lines are given by a Hamiltonian $\psi_{p}(\psi_{t},\theta,\varphi,t)$ in canonical form with time $t$ just a parameter:
@@
In an axisymmetric torus, the magnetic surfaces can be bounded by X-points, as is well known because of tokamak divertors. At an X-point, the gradient of the magnetic poloidal angle vanishes, $\vec{\nabla}\theta=0$, which implies $\vec{B}\cdot\vec{\nabla}\theta=$ and requires $\vec{\nabla}\psi_{t}\rightarrow\infty$ since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi\neq 0$.
The shortest distance between neighboring irrational magnetic surfaces $\Delta(\theta,\varphi)\equiv|\psi_{t}|/|\vec{\nabla}\psi_{t}|$ can have an exponentially large ratio $\Delta_{max}/\Delta_{min}=e^{2\sigma}$ despite all the surfaces remaining perfect. As discussed in Section VI, when $\sigma\gtrsim 20$ even the small resistivity in fusion plasmas will cause the magnetic surfaces to break and form a region of volume-filling chaotic magnetic field lines. Since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$ varies little over a magnetic surface compared to $e^{20}$, places where neighboring surfaces are close together must be places where the field lines in the surfaces are far apart and vice versa (Boozer, 2022).
III Breakup of magnetic surfaces
Magnetic surfaces breakup by the formation of magnetic island when a rational surface $\iota=n/m$ is perturbed with the width of the island proportional to the square root of the perturbation. The interior of an island has magnetic surfaces. The magnetic surfaces in a curl-free stellarator field can be considered to be those of an island. Figure 2 shows islands can have islands in their interiors.

### v2

@@ 1 Introduction
is neither zero nor infinity in the spatial region in which they are to be used. Defining the position function just means giving $x$, $y$, and $z$ as functions of $(\psi,\theta,\varphi)$. The appendix to (Boozer:2004) explains how arbitrary coordinate systems, such as $\vec{x}(\psi,\theta,\varphi)$, are used in calculations.
@@
One reason coordinate systems are important to a discussion of the properties of magnetic field lines, is their trajectories are given by a Hamiltonian of the $H(p,q,t)$ form. The Hamiltonian, its canonical momentum $p$ and canonical coordinate $q$, and its canonical time $t$ have a subtle relationship to physical quantities. Section 2 shows the Hamiltonian is the poloidal flux $\psi_{p}(\psi_{t},\theta,\varphi,t)$, the canonical momentum is the toroidal magnetic flux $\psi_{t}$, the canonical coordinate is the poloidal angle $\theta$, and the canonical time the toroidal angle $\varphi$, Figure 1. The actual clock time $t$ is just a parameter in the Hamiltonian.
@@
(Boozer:1983) showed the topological properties of magnetic field lines at each instant in time are given by $\psi_{p}(\psi_{t},\theta,\varphi,t)$. The position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ is needed to plot the field lines in ordinary space with $\partial\vec{x}/\partial t$ giving the velocity of $(\psi_{t},\theta,\varphi)$ through space, which is the field line velocity when field line topology is not changing, which implies $\psi_{p}$ can be taken to be independent of time.
@@
When a particular line $\vec{x}_{0}(\ell)$ remains within a bounded region of space, it has three possibilities as $\ell\rightarrow\infty$: It can close on itself after a distance $\ell_{0}$. It can cover a surface coming arbitrarily close to every point on a surface without ever going through the same point twice. It can cover a volume coming arbitrarily close to every point in a volume of space without ever going through the same point twice. All three types of trajectories are illustrated by the Standard Map, Figure 2. The iteration number of the map represents the distance $\ell$. The periodicity of both coordinates of the Standard Map makes it non-trivial to translate them into the $(\psi_{t},\theta)$ coordinates of the magnetic field line Hamiltonian $\psi_{p}$. (Boozer-Punjabi:2018) in their Appendix B have obtained an explicit formula for a $\psi_{p}(\psi_{t},\theta,\varphi)$ that represents the qualitative features of stellarator equilibria including the non-resonant divertor that (Strumberger:1992) found naturally arises at the edge optimized stellarators. Non-resonant divertors are discussed in Section 5.2 as way to build intuitive understanding of the subtleties of cantori and turnstiles.
@@
Figure 2: The Standard Map gives trajectories by iteration in a space of two periodic coordinates. The horizontal axis is $\alpha$, the vertical axis is $\beta$, and both have a period of unity. The $n+1$ iterate of the map is $\alpha_{n+1}=\alpha_{n}+\beta_{n+1}$ and $\beta_{n+1}=\beta_{n}-(k/2\pi)\sin(2\pi\alpha_{n})$. Trajectories are illustrated for $k=0.975.$ Black trajectories show island chains. The red region is a single trajectory iterated $10^{6}$ times, but the trajectory still escapes from that region after approximately $10^{9}$ iterations. This was Figure 1 in J. D. Meiss, Thirty years of turnstiles and transport, Chaos 25, 097602 (2015).
@@
The behavior of the magnetic field lines that are infinitesimally separated from an arbitrary line $\vec{x}_{0}(\ell)$ can be studied without approximation by a particularly simple Hamiltonian (Boozer:2012). The study of neighboring lines, which is reviewed in Appendix A, showed: (1) It is possible for field lines infinitesimally separated from a given line $\vec{x}_{0}(\ell)$, of any of the three types, to have separations that increase or decrease exponentially with $\ell$. There must be an equal flux of neighboring lines that diverge from and approach the line $\vec{x}_{0}(\ell)$ due to $\vec{\nabla}\cdot\vec{B}=0$. When lines that lie on a circle of infinitesimal radius at $\ell=0$ are followed a distance $\ell$, the circle will deform into an ellipse. The field strength times the square of the semi-major axis of the ellipse increases as $e^{\sigma_{max}(\ell)}$, and the field strength times the square semi-minor axis decreases as $e^{-\sigma_{max}(\ell)}$. (2) When the magnetic field depends on only two spatial coordinates, a single field line cannot cover a volume, as in the red area of Figure 2. Three spatial coordinates are required. Two-coordinate magnetic fields, as in axisymmetry, can have X-points (more correctly X-lines) which are specific field lines from which neighboring lines exponentially approach and separate. However, the volume in which lines have at least a certain exponentiation is itself exponentially small. (3) Curl-free magnetic fields can exhibit field lines with an exponential separation throughout a volume. An ideal magnetic evolution does not conserve the exponentiation $\sigma_{max}(\ell)$. In other words, the exponentiation can be a property of the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ with $\psi_{p}$ an unchanging function of its canonical coordinates. The current density required for an ideal evolution to produce an exponentiation $\sigma_{max}$ is proportional to $\sigma_{max}$, Appendix A.3. As will be discussed in Section 6, the time required for a large scale reconnection to occur is determined by a timescale set by the ideal evolution times $\sigma_{r}$. The required exponentiation $\sigma_{r}$ for large scale reconnection is the ratio of timescale for non-ideal diffusive effects, such as resistivity, to a timescale determined by the ideal evolution. As discussed in Section 6.2, the time scale for reconnection is determined by an ideal timescale times the natural logarithm of $\sigma_{r}$.
@@
The tendency to construct two dimensional models has led to the prejudice that exponential separation of magnetic field lines throughout a volume is peculiar. Actually any magnetic field that depends on all three spatial coordinates tends to be chaotic, which is illustrated by the design of a curl-free magnetic fields for a stellarators. Obtaining nested magnetic surfaces, as required for a stellarator, rather than having the field lines fill a volume requires careful optimization. Even a small perturbation to an optimized curl-free stellarator can destroy the surfaces. Nevertheless, axisymmetric tokamaks are even more sensitive to external field perturbations due to their amplification by coupling to marginally stable current-driven kinks (Park:2010).
@@
The concept of magnetic field line chaos and its definition are sufficiently subtle that Section 4 is largely devoted to the topic.
@@
When the magnetic field lines lie in toroidal magnetic surfaces, the poloidal angle $\theta$ can be chosen so the poloidal flux is a function of only the toroidal flux and time, $\psi_{p}(\psi_{t},t)$. At each instant in time, the trajectories of field lines are then given by $\psi_{t}$ and $\theta_{0}$ constant with $\theta=\theta_{0}+\iota(\psi_{t})\varphi$, where the rotational transform $\iota\equiv\partial\psi_{p}/\partial\psi_{t}$. In tokamaks, it is customary to use the safety factor $q\equiv 1/\iota$ instead of the rotational transform.
@@
On a rational surface the rotational transform is a rational number, the ratio of two integers $\iota=N/M$. The field lines on that surface close after $N$ poloidal transits and $M$ toroidal transits. This identification with poloidal and toroidal transits may seem backwards from the standard convention of using $m$ as a poloidal and $n$ as a toroidal mode number as in $\cos(m\theta-n\varphi)$. But, when $\iota=N/M$, then along a field line $\cos(m\theta-n\varphi)=\cos\Big(m\theta_{0}+(mN/M)\varphi-n\varphi\Big)$, which has no variation along the line since it equals $\cos(m\theta_{0})$ when $m=M$ and $n=N$.
@@
(Boozer:2004) showed the time derivative of the poloidal flux is the loop voltage,
$\frac{\partial\psi_{p}(\psi_{t},t)}{\partial t}=V_{\ell}.$ (2)
At the magnetic axis, Equation (2) is a trivial consequence of Faraday’s Law plus Stokes’ Theorem with the loop voltage $V_{\ell}$ the integral of the electric field around the axis. The proof of this equation for an arbitrary magnetic surface, which is given in the Appendix to (Boozer:2004), is non-trivial. On an arbitrary magnetic surface, the loop voltage is an integral along a field line at an instant of time,
@@
As proven in the paragraph below Equation (17), when $\vec{\nabla}V_{\ell}=0$ the topology of the magnetic field lines cannot change, and their evolution is considered ideal. This is in contrast to ideal magnetohydrodynamics, which requires a zero resistivity $\eta$, and that implies $V_{\ell}=0$. As pointed out by (Boozer:2025:Constraints) and discussed in Section 7.1, the distinction between $\vec{\nabla}V_{\ell}=0$ and $V_{\ell}=0$ is important when devising strategies for avoiding disruptions in tokamaks. The condition $\vec{\nabla}V_{\ell}=0$ is consistent with changing the magnetic flux at the axis of a toroidal plasma without changing the magnetic field in the plasma while the condition $V_{\ell}=0$ is not.
@@
Section 3 explains a perturbation that breaks magnetic surfaces and forms chains of magnetic islands at the rational surfaces that resonate with the rotational transform $\iota\equiv\partial\psi_{p}/\partial\psi_{t}$. A resonant perturbation to $\psi_{p}$ has a toroidal mode number $n$ and a poloidal mode number $m$ with $n/m=\iota$. The width of these islands is proportional to the square root of the resonant perturbations. Non-resonant perturbations distort magnetic surfaces but do not break them. Resonant perturbations that create islands cannot be ideal, they must change the form of the poloidal flux, but the position function can be held fixed.
@@
Non-resonant perturbations can be taken to change the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$ while the poloidal flux can be held fixed using a canonical transformation or what is mathematically equivalent an electric potential. This is shown in the appendix to (Boozer:2004).
Section 4 discusses the subtle properties of chaotic magnetic field lines. These properties need to be understood before undertaking a study of cantori and turnstiles, which serve as boundaries on region of chaotic magnetic field lines.
@@
(Chirikov:1960) gave the approximate condition for the creation of large chaotic volumes: the overlap of the islands from different resonant rational surfaces.
As discussed in Section 5, the actual formation of chaotic regions is more complicated than just the overlap of island chains. (MacKay:1984) showed that chaotic regions are formed by cantori and the turnstiles that penetrate them.
@@
The last magnetic surface to break between two large magnetic island chains is the most irrational surface between the resonant surfaces associated with those chains, Section 5. This is quantified by the magnitude of $m$ required for the transform on this surface to satisfy $\iota(\psi_{t})=n/m+\epsilon_{r}$ with $n$ a toroidal mode number and $m$ a poloidal mode number as $|\epsilon_{r}|\rightarrow 0$. As the last magnetic surface breaks, it is called a cantorus. A cantorus is like an irrational magnetic surface but develops holes. These holes are places where $\vec{B}\cdot\hat{n}\neq 0$ with $\hat{n}$ the unit normal to the cantorus. The holes define highly collimated tubes of magnetic flux and must come as an inward and an outward pair due to $\vec{\nabla}\cdot\vec{B}=0$. For this reason they are called turnstiles.
@@
The highly collimated flux tubes of turnstiles are responsible for the severe damage that can be caused by runaway electrons following the lines in these tubes to localized positions on the chamber walls (Boozer-Punjabi:2016). This collimation also responsible for the formation of a natural non-resonant divertor at the edge of a stellarator (Strumberger:1992) and (Boozer-Punjabi:2018). Section 5.2 on non-resonant divertors provides an intuitive introduction to cantori and turntiles through an application to a practical problem.
If the plasma evolution were perfectly ideal, a resonant perturbation would affect only the position function $\vec{x}(\psi_{t},\theta,\varphi,t)$. But, a resonant ideal perturbation forces the magnetic field to have a delta-function current density at the resonant rational surface $\iota=n/m$. This current causes distortions to the neighboring surfaces producing magnetic perturbations with all harmonics $k$ of the resonant perturbation $(kn)/(km)=\iota$ with $k$ any integer (Huang-Hudson:2022). In a torus, the toroidicity produces magnetic perturbations with poloidal harmonics of strength $\epsilon^{k}$ at $m\pm k$ with $\epsilon$ the inverse aspect ratio of the torus. Any rational number can be produced by the harmonics of $n/m$ together with the harmonics of $m$ produced by toroidicity. The more ideal the perturbation, the more contorted the magnetic surfaces become, which leads to their rapid breaking by an arbitrarily small non-ideal effect, such as resistivity (Boozer:2022). This is discussed in Section 6.
@@
Section 7 gives additional examples of the application if the concepts of magnetic field line chaos, cantori, and turnstiles to toroidal fusion plasmas but also expresses concern about the future of innovation within U. S. fusion program.
2 Representation of a magnetic field and its field lines
@@
(Boozer:1983) showed that the magnetic field in a toroidal region can always be written as
@@
Equation (4) follows from writing the vector potential $\vec{A}$, where $\vec{B}=\vec{\nabla}\times\vec{A}$, using three coordinates $(r,\theta,\varphi)$. Then, $\vec{A}=A_{r}\vec{\nabla}r+A_{\theta}\vec{\nabla}\theta+A_{r}\vec{\nabla}\varphi+\vec{\nabla}g$. The gauge can be chosen so $\partial g(r,\theta,\varphi)/\partial r=-A_{r}$. Let $\psi_{t}/2\pi\equiv A_{\theta}+\partial g/\partial\theta$, and $\psi_{p}/2\pi\equiv-(A_{\varphi}+\partial g/\partial\varphi)$, so $2\pi\vec{A}=\psi_{t}\vec{\nabla}\theta-\psi_{p}\vec{\nabla}\varphi$, which has Equation (4) as its curl.
In general the poloidal flux depends on all three spatial coordinates and time, $\psi_{p}(\psi_{t},\theta,\varphi,t)$. (Boozer:1983) showed the trajectories of the magnetic field lines are given by a Hamiltonian $\psi_{p}(\psi_{t},\theta,\varphi,t)$ in canonical form with time $t$ just a parameter:
@@
In an axisymmetric torus, the magnetic surfaces can be bounded by X-points, as is well known because of tokamak divertors. At an X-point, the gradient of the magnetic poloidal angle vanishes, $\vec{\nabla}\theta=0$, which implies $\vec{B}\cdot\vec{\nabla}\theta=0$ and requires $\vec{\nabla}\psi_{t}\rightarrow\infty$ since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi\neq 0$.
The distance between neighboring irrational magnetic surfaces $\Delta(\theta,\varphi)\equiv|\psi_{t}|/|\vec{\nabla}\psi_{t}|$ can have an exponentially large ratio between its maximum and minimun, $\Delta_{max}/\Delta_{min}=e^{2\sigma}$, despite all the surfaces remaining perfect. As discussed in Section 6, when $\sigma\gtrsim 20$ even the small resistivity in fusion plasmas will cause the magnetic surfaces to break and form a region of volume-filling chaotic magnetic field lines. Since $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$ varies little over a magnetic surface compared to $e^{20}$, places where neighboring surfaces are close together must be places where the field lines in the surfaces are far apart and vice versa (Boozer:2022).
3 Breakup of magnetic surfaces
Magnetic surfaces breakup by the formation of magnetic islands when a rational surface $\iota=n/m$ is perturbed with the width of the island proportional to the square root of the perturbation. The interior of an island has magnetic surfaces. The magnetic surfaces in a curl-free stellarator field can be considered to be those of an island. Figure 2 shows islands can have islands in their interiors.